Density of 3M solution of NaCl is 1.252g ml^-1 the molality of solution is ?

Grade:11

1 Answers

naimatullah khan
5 years ago
molarity=no. of moles of solute /volume of solutionn
given molarity =3M
means, no of moles of solute =3 and volume of solution =1L
density =mass /volume 
1.252=mass/1000                       (1L=1000ml)
mass of solution =1252g
no. of moles of solute = mass (W)/molar mass(M)
3=W/58.5                                                                    (NaCl= 23+35.5=58.5g/mol)
W=175.5g
mass of solvent =mass of solution-mass of solute 
mass of solvent =1252-175.5=1076.5g
molality =no. of mole of solute /mass of solvent in kg 
molality =3/1.076
            =2.788